Grade 1010 grade maths

A sum of Rs 25000 is deposited in a co operative society for 2 years. The rate of compound interest is 8% pa for the first year and 10% pa for the second year where the interest is added half yearly in both cases. What will be the compound amount and compound interest on the sum?

TheMiralce123
6 Years agoGrade 10

1 Answer

Saurabh Koranglekar
6 Years ago
Dear student

Compound amount is (25000*(1.04)^ 2) for first yr = 27040

Compound amount at end of second yr = 27040 * (1.05)^2 = 29811.6

Compounded interest is 4811.6

Regards